the perimeter of a rectangle is 700 yards. what are the dimensions of the rectangle if the length is 70 yards more then the width
Answer by JulietG(1812)

The perimeter is the fence around. It requires 2 pieces of length & 2 pieces of width. P = 2L + 2W
The question gives the following information:
700 = 2L + 2W
L = 70 + W
.
We can plug in the known values and then solve for W
700 = 2L + 2W
700 = 2(70+W) + 2W
700 = 140+2W + 2W
700 = 140 + 4W
700-140 = 4W
560 = 4W
560/4 = W
140 = W
.
If the width is 140 and the length is 70 more than that, then the length is 140+70 = 210
Let's make certain by adding these
2(140) + 2(210) = 700
280 + 420 does indeed = 700
Success!